OpenStudy (dioneches):

Convert 62 degrees celcius to fahrenheit Please help will give medal

OpenStudy (anonymous):

Have you googled "Celsius to Fahrenheit conversion equation"? Seems like it would be common sense to do that `Celsius to Fahrenheit`\[\huge F= \frac{9}{5}C+32\] `Fahrenheit to Cesius`\[\huge C=(F-32)\cdot \frac{5}{9}\]

OpenStudy (dioneches):

sorry... I was on here and i thought i should ask... I suck at this stuff...

OpenStudy (anonymous):

X) In the future it might save time to see if you can look it up first and then ask if you can't find anything. It usually takes a while to get a response in the physics section

OpenStudy (dioneches):

ok... sorry...

OpenStudy (anonymous):

Nothing to be sorry about! X) I'm just telling you it might be quicker to Google something that simple given that the physics section can be really dead a lot of times

OpenStudy (dioneches):

ok...